when the first of two consecutive positive integers is added to the square of the second the result is five.Find the two integers
Answer by Maths68(1474)

Let
one integer = x
other integer = x+1
===============================================
Given
When one added to the square of the second the result is five
x+(x+1)^2=5
x+x^2+2x+1=5
x+x^2+2x+1=5
3x+x^2=5-1
x^2+3x=4
x^2+3x-4=0
x^2+4x-x-4=0
x(x+4)-1(x+4)=0
(x-1)(x+4)=0
x-1=0 or x+4=0
x=1 or x=-4 (We need positive integer so -4 is not valid)
one integer = x = 1
other integer = x+1 = 1+1 = 2
=============================================================
Check
first + square of the other = 5
1+2^2=5
1+4=5
5=5